# Diffusive Spreading Across Dynamic Mitochondrial Network Architectures

### v1

@@ Abstract
Networks of physical units can vary from a stationary set of spatially-embedded links to a collection of mobile agents that undergo transient social interactions. In living cells, mitochondria form architectures that span across these regimes, transitioning between fragmented, partly connected, and highly fused structures depending on cell type and state. Diffusive transport of biomolecular components through these networks helps to homogenize the mitochondrial population. Here we address the connection between dynamic network architecture and the rate of diffusive mixing through simulations and analytic models that incorporate fusion, fission, and rearrangement. We find that the material delivered from a source to the rest of the network depends on the network dimensionality and a balance of competing timescales for encounter, fusion, and diffusive dispersion. These results provide a quantitative basis for predicting the homogenization of proteins, lipids, ions, or genetic material through the mitochondrial population. The general principles identified in this work capture diffusive spreading through both social and physical networks, unifying a continuum of spatial network architectures.
@@ Introduction
Diffusive transport through networks has been studied in a variety of contexts, including disease spread in epidemiology [1, 2], innovations in social networks [3], communication among insects [4, 5], nutrient flows in fungal networks [6], diffusion of signals and nutrients in the brain extracellular space [7, 8], and oil recovery in porous rock [9]. These systems can be separated into two broad classes: one where the topology of network connections is stationary and limited to nearest neighbors, and another where network nodes are mobile and promiscuous, interacting with many different partners over time.
Physical networks form an important category of stationary, spatially-constrained network structures [10]. In such networks, edges represent objects subject to physical limitations (space-filling, steric repulsion, etc.) [11, 12] connecting degree-1 tips and degree-3 junctions, with higher degree nodes exceedingly rare [13, 14]. Examples include fungal mycelia [15], porous rocks [16], and neuronal synaptic networks [17]. Transport behaviors on stationary networks can be described by the graph Laplacian, whose eigenvalues govern the spreading timescales [18, 19, 20]. In some percolation problems, physical networks become dynamic as edges are allowed to flicker between an active and inactive state [21, 22].
@@
Despite the distinct modeling approaches employed, physical and social networks lie on a continuum of temporal network structures [23] with varying timescales of topological rearrangement. The intracellular environment exemplifies this spectrum. The endoplasmic reticulum forms a highly-looped physical lattice of membrane-bound tubules that enables diffusive transport of proteins [25] and ions [26] across the entire cell. At the opposite extreme, the population of endocytic vesicles constitutes a social network of discrete transiently-interacting compartments [27, 28].
Mitochondria form another intracellular network with striking structural variability. Mitochondrial networks transition between fragmented and hyperfused architectures in response to disease state [29, 30, 31, 32, 33, 34], metabolic conditions [35, 36, 37], calcium signalling and apoptosis [38, 39], cell division [40, 41] and cell type [42, 43, 44, 45]. These structural changes are thought to support important functions such as genetic complementation [46, 47], dilution of harmful reactive oxygen species (ROS) [48], power cabling [49, 50], quality control [51, 52, 53, 54], and modulation of mitochondrial heterogeneity across the population [55, 56]. These cellular functions rely on the connectivity-dependent diffusive spread of proteins, lipids, and/or ions through the underlying mitochondrial network [57, 22].
To elucidate how mitochondrial structure and dynamics govern diffusive mixing, we introduce a quantitative framework for transport on spatially-embedded dynamic networks. We show that network connectivity governs spreading behavior by modulating several competing timescales: the cluster filling rate, encounter rate, and fusion rate, as well as the decay rate of the spreading material. In particular, we demonstrate a transition from rapid but low-dimensional spread in well-connected networks to slower three-dimensional spreading through an interacting population. Due to the different dimensionalities involved, particles with distinct diffusivities and networks with distinct structures can exhibit qualitatively different scaling behavior for dispersion from a source. The network architectures considered here are motivated by observations of three-dimensional mammalian mitochondria, whose structural diversity exemplifies a broad range of spatial networks.
@@
We seek to quantify the rate of material dispersion through dynamic spatial networks ranging from the social to the physical regime. These limiting architectures, as well as a continuum of intermediate network connectivities, are encompassed by a simulation framework [58], illustrated in Fig. 1, consisting of $N_{0}$ interacting spherocylindrical units of length $\ell_{0}$, which move diffusively through a bounded spherical domain of radius $R$ with unit diffusivity $D_{1}$. Upon close encounters, pairs of units may fuse into larger clusters with rate constants $k_{u1}$ (for tip-tip fusions) or $k_{u2}$ (for tip-side fusions). Fission breaks connections at the nodes, with fission rate proportional to the number of attached edges ($k_{f}$ at degree-2 nodes and $1.5k_{f}$ at degree-3 nodes), in keeping with prior models [14]. At steady state, the system exhibits a characteristic mean cluster size $\langle n\rangle$ set by the balance of fusion and fission, and an associated cluster diffusivity $D_{n}$. Details of this structural model and its relevant parameterization for mitochondrial networks in different cell types are described in Ref. [58]; here we focus on the spreading of material throughout the resulting dynamic networks.
@@
We select at random a single unit, which serves as a particle source with fixed constant concentration $h_{0}=1/\ell_{0}$. Such a fixed source could represent, for instance, a mitochondrion that buffers calcium at a contact site with the endoplasmic reticulum [59], or one that produces a regulated amount of mitochondrially encoded The concentration field is propagated between connected units on the network with diffusivity $D_{p}$, using a finite volume approach [60] discretized at the level of individual units. Concentrations decay with a constant rate $k_{d}$ (representing removal from the network). The decay rate serves to set a relevant timescale, with the resulting calculations closely related to the question of how much material spreads from the source over a certain time. We quantify dispersion by computing the total amount of material ($S$) in the network at steady-state, excluding the source unit. Example steady-state snapshots are shown in Fig. 1b and Supplemental Video 1.
@@
Given the complex internal structure of mitochondria and the variety of functionally relevant biomolecules within them, particle diffusivities can vary broadly. For example, proteins in the inner mitochondrial membrane can become trapped in the extensive folds of cristae [61]. The cristae also serve as diffusive barriers to hinder transport of solutes in the matrix [62] and intermembrane space [63]. Our simplified modeling approach coarse-grains these complications into a single effective diffusivity $D_{p}$ representing the motion of a particle along the mitochondrial tubule axis. Reported effective diffusivities vary over orders of magnitude, from $\sim 0.004\mu\text{m}^{2}/\text{s}$ for ATP synthase components  [64], to $\sim 20\mu\text{m}^{2}/\text{s}$ for mitochondrial matrix proteins [65], with diffusivity of ions presumed even higher [66]. We consider a range of diffusivities $D_{p}=0.4-40\mu\text{m}^{2}/\text{s}$ in our simulations. The diffusivity of individual mitochondria is estimated to be much slower, at $D_{1}\approx 0.25\mu\text{m}^{2}/\text{min}$ [42, 67].
@@
The results below are reported in dimensionless units, relative to the length scale $2\ell_{0}$ and the timescale $1/k_{f}$. A reasonable estimate sets the mitochondrial unit length to $\ell_{0}=0.5\mu$m and the fission timescale to $1/k_{f}\approx 2\text{ min}$ [58], consistent with experimental measurements of overall fission rate [68, 42]. The corresponding dimensionless diffusivities are $D_{p}=48-4800$ for the material, and $D_{1}=0.5$ for individual mitochondria.
@@ Results
In the limit where $\tau_{d}$ is the shortest timescale, material spreads primarily within a single connected component. This regime is relevant for highly fused networks with large cluster sizes. Such networks also tend to be relatively static in their topology, with any fissions that occur rapidly followed by re-fusion with the same neighboring unit [58]. For static networks, the steady-state distribution of material can be found by solving the diffusion equation on each one-dimensional edge, while matching boundary conditions at each node (details in SI Appendix). This method provides an exact solution for each individual network structure. However, to gain insight on the relevant scaling regimes, we turn to a mean-field continuum approach.
Figure 2: Material spreading on static networks of different dimension $d$. Two different sets of networks are considered: near-linear structures (blue) with $k_{u2}=0.01k_{u1}$ and highly-branched structures (green) with $k_{u2}=10k_{u1}$. The simulations are run to steady state, and the network structures are then frozen. (a) Calculation of the network dimension for a single instance of each network type (shown in inset). The number of nodes within a given graph distance is plotted against the graph distance on log-log axes, with the slope giving the dimension, $d$. Averaging over 21 snapshots from 3 independent simulations yields effective dimensionalities $d=1.2,d=1.9$, respectively. (b) Total material delivered from a source unit on the static networks is plotted as a function of the diffusive lengthscale, $\lambda_{p}=\sqrt{D_{p}/k_{d}}$. Solid curves show exact solution (details in SI Appendix), averaged over replicate snapshots. Dashed-circle curves show approximation with the linear motif (inset i), applicable at short $\lambda$. Dashed-square curves show continuum solution on a fractal domain (inset ii), matching well at long $\lambda$.
@@
For the regime where material spreads over a more extensive network structure ($\lambda_{p}$ much greater than segment length), we consider diffusion through a $d$-dimensional medium. Here we approximate the spreading as spherically symmetric, with a $d-$sphere of radius $a_{n}$ (enclosing the source edge) maintained at fixed concentration, and a reflecting $d-$sphere of radius $R_{n}$ representing the outer boundary of the cluster. We set $R_{n}$ such that the average distance between two points in the $d$-dimensional domain is equal to the average graph distance between network nodes (details in SI Appendix). Assuming that the $n\right>$ network units are uniformly distributed within the continuum sphere, the radius $a_{n}$ is set proportionately to allow for 1 source unit within the inner sphere: $1/a_{n}^{d}=(\langle n\rangle-1)/(R_{n}^{d}-a_{n}^{d})$.
$\begin{gathered}S_{d}=\frac{d}{x_{1}}\frac{\left[-I_{\nu-1}(x_{1})K_{\nu-1}(x_{2})+I_{\nu-1}(x_{2})K_{\nu-1}(x_{1})\right]}{\left[I_{\nu}(x_{1})K_{\nu-1}(x_{2})+I_{\nu-1}(x_{2})K_{\nu}(x_{1})\right]}\\ \end{gathered}$ (1)
Fig. 2 shows the correspondence between diffusion through a fractal continuum (dashed-square curves) and the exact solution (solid curves) for networks with different connectivities. The linear motif solution $S_{1}$ (dashed-circle curves) is a good estimate for small diffusive lengths, while the continuum solution $S_{d}$ (dashed-square curves) is a better approximation for large $\lambda_{p}$. When the diffusive lengthscale is large enough ($\lambda_{p}>R_{n}$), the entire connected component saturates. Overall, material delivery is boosted by increased network branching [22], which increases both the network dimension and the network density (lower $R_{n}$ for the same number of network units).
@@
We next consider the regime of highly fragmented ‘social’ networks, where clusters are sufficiently small (and particle diffusivity sufficiently fast) that concentrations are fully equilibrated within each cluster: $\tau_{c}\ll\left\{\tau_{\text{d}},\tau_{\text{int}}\right\}$. In this limit, we can approximate the network as a system of $N=N_{0}/\left<n\right>$ identical spherical units with effective radius $a$, each representing a cluster of uniform concentration. These effective units are capable of undergoing transient fusions whenever they are within contact radius $b$, with rate constant $k_{u}$. Each fusion equilibrates the particle concentration in the two units involved, and is instantaneously followed by fission so that no larger structures are formed.
@@
$\frac{dh_{c}(t)}{dt}$ $=k_{u}(h_{0}\ell_{0}-h_{c}(t))-\frac{I}{\rho v_{c}}-k_{d}h_{c}(t),$ (2b)
where $k_{u}(h_{0}\ell_{0}-h_{c}(t))$ represents the injection of new material into the system via fusion with the source, $I=-4\pi b^{2}\rho D\left.\frac{dh(r,t)}{dr}\right|_{b}$ is the current of material leaving the contact zone, $\rho=(N_{0}-\langle n\rangle)/V$ is the density of network units, $V=4/3\pi(R^{3}-a^{3})$ is the domain volume, and $v_{c}=4/3\pi(b^{3}-a^{3})$ is the contact zone volume. Interactions between non-source units do not alter the mean-field concentrations, and the overall material in the system can only increase upon encounters with the source.
Figure 3: Material delivered from a source unit in a system of fragmented clusters. Solid lines show mean-field solutions (see Eq. 3) with unit cluster size $\langle n\rangle=1$, for different fusion rates $k_{u}$. Colored dots show explicit simulation results for a simplified system of interacting spheres with uniform size (inset). Dashed black lines show the limiting behavior for fusion-limited and diffusion-limited regimes.
We compute the steady-state solution of Eq. 2 and integrate $h(r)$ over the space to find the total amount of material ($S$) in the system, excluding the source unit (details in SI Appendix):
@@
The total material in the system exhibits distinct scaling behaviors (Fig. 3, dashed) depending on the relative timescales for decay, encounter ($\tau_{\text{enc}}=(4\pi Db\rho)^{-1}$), and fusion ($\tau_{u}=(k_{u}\rho v_{c})^{-1}$). In keeping with standard results for diffusion to a partially reactive target [71], the overall interaction time is given by the sum of waiting times for the two-step process: $\tau_{\text{int}}=\tau_{\text{enc}}+\tau_{u}$. In the limit of fast decay ($\tau_{d}\ll\tau_{\text{int}}$), the steady-state material in a social network of unit size clusters approaches $S\rightarrow\rho v_{c}h_{c}=\rho v_{c}k_{u}/k_{d}$. In this limit, the total amount delivered is set by the number of units in the contact zone ($\rho v_{c}$) and how often they fuse with the source during the decay time ($k_{u}/k_{d}$), with no spatial spread.
@@
In a third limit, fusion of clusters is fast and decay is slow relative to the timescale of diffusive encounter. ($\tau_{u}\ll\tau_{\text{enc}}\ll\tau_{d}$). For this regime, the total material content is given by $S\rightarrow\rho zh_{c}\rightarrow 4\pi Db\rho/k_{d}$. This is a diffusion-limited regime, where clusters arriving at the contact region fuse with the source nearly instantaneously ($h_{c}\approx h_{0}\ell_{0}=1$), and spreading is determined by the balance of arrival and decay rates.
Fig. 3 shows the full solution of the mean-field model (Eq. 3) for spreading on a fragmented network, with variable fusion rate $k_{u}$ and a unit cluster size $n\right>=1$. The analytic calculations accurately reproduce simulations with diffusive spheres that exchange material via transient fusion (dots in Fig. 3, details in SI Appendix).
Figure 4: Spreading on simulated networks exhibits a transition between the static network and social network regime. The total material is plotted as a function of the decay time for different fusion rate constants $k_{u1}$ (solid lines). Approximation of the network as a fractal continuum captures the behavior in the large cluster, fast decay limit, while the mean-field social network approximation describes the small cluster, slow decay limit. The combined solution (Eq. 4, dashed curves) encompasses the full range of decay timescales and particle diffusivities. The transition between the two approximations occurs when the cluster containing the source unit is filled ($S=\langle n\rangle-1$), as marked by colored arrows in (a). Insets show example simulation snapshots for the parameters indicated by the gray lines. Parameters $N_{0}=250,\ell_{0}=0.5,R=5,k_{u2}=\frac{1}{3}k_{u1},k_{f}=1$ are used in both (a) and (b), with the particle diffusivity reduced from $D_{p}=4800$ in (a) to $D_{p}=48$ in (b).
@@
The model for spreading through a social network of interacting fragments can be expanded to approximate a regime with larger clusters. We use simulations of dynamic networks with different structures [58] to extract the effective parameters (mean cluster size $n\right>$, effective unit diffusivity $D_{n}$, steric radius $a_{n}$, contact volume $v_{c}$, and effective local fusion rate $k_{u}$) for the simplified social network model (details in SI Appendix).
When individual clusters are homogeneously filled upon each interaction ($\tau_{c}<\left\{\tau_{d},\tau_{\text{int}},\tau_{f}\right\}$), the solution in Eq. 3 can be used directly with the appropriate values of the parameters $a,v_{c},k_{u}$. However, for slowly diffusing particles or very large clusters ($\tau_{c}>\tau_{d}$), the source cluster may be only partly full of material. The average concentration in the source cluster, $h_{0}^{(d)}$, is then determined by the static solution (Eq. 1) within the cluster: $\ell_{0}h_{0}^{(d)}=(S_{d}(\sqrt{D_{p}/k_{d}})+1)/\langle n\rangle$. When another cluster fuses with the source cluster, fission may terminate the encounter before filling is complete ($\tau_{f}<\tau_{c}$). The amount transferred should then be scaled by the fraction of the new cluster that is filled prior to fission: $f=(S_{d}(\sqrt{D_{p}/k_{f}})+1)/\langle n\rangle$. These two corrections modify the dynamic Eq. 2 for the concentration in the contact zone and its solution $S$ as follows:

### v2

@@ Abstract
Networks of physical units can vary from a stationary set of spatially-embedded links to a collection of mobile agents that undergo transient social interactions. In living cells, mitochondria form architectures that span across these regimes, transitioning between fragmented, partly connected, and highly fused structures depending on cell type and state. Diffusive transport of biomolecular components through these networks limits the heterogeneity of the mitochondrial population. Here we address the connection between dynamic network architecture and the rate of diffusive mixing through simulations and analytic models that incorporate fusion, fission, and rearrangement. We find that the material delivered from a source to the rest of the network depends on the network dimensionality and a balance of competing timescales for encounter, fusion, and diffusive dispersion. We extract morphological and dynamic parameters for mitochondrial networks in three human cell lines, demonstrating that different cells span across both the physical and social network regimes. Mixing in mitochondrial networks is shown to be limited by material transport through connected tubules for slowly-diffusing particles and by inter-mitochondrial encounter rates for rapidly diffusing ones. These results provide a quantitative basis for predicting the homogenization of proteins, lipids, ions, or genetic material through the mitochondrial population. The general principles identified in this work capture diffusive spreading through both social and physical networks, unifying a continuum of spatial network architectures.
@@ Introduction
Diffusive transport through networks has been studied in a variety of contexts, including disease spread in epidemiology [1, 2], innovations in social networks [3], communication among insects [4, 5], diffusion of signals and nutrients in the brain extracellular space [6, 7], and oil recovery in porous rock [8]. These systems can be separated into two broad classes: one where the topology of network connections is stationary and limited to nearest neighbors, and another where network nodes are mobile and promiscuous, interacting with many different partners over time.
Physical networks form an important category of stationary, spatially-constrained network structures [9]. In such networks, edges represent objects subject to physical limitations (space-filling, steric repulsion, etc.) [10, 11] connecting degree-1 tips and degree-3 junctions, with higher degree nodes exceedingly rare [12, 13]. Examples include fungal mycelia [14], neuronal synaptic networks [15], and interconnected pores in rocks [16]. Transport behaviors on stationary networks can be described by the graph Laplacian, whose eigenvalues govern the spreading timescales [17, 18, 19]. In some percolation problems, physical networks become dynamic as edges are allowed to flicker between an active and inactive state [20, 21].
@@
Despite the distinct modeling approaches employed, physical and social networks lie on a continuum of temporal network structures [22] with varying timescales of topological rearrangement. The intracellular environment exemplifies this spectrum. The endoplasmic reticulum forms a well-connected physical lattice of membrane-bound tubules that enables diffusive transport of proteins [24] and ions [25] across the entire cell. At the opposite extreme, the population of endocytic vesicles constitutes a social network of discrete transiently-interacting compartments [26, 27].
Mitochondria form another intracellular network with striking structural variability. Mitochondrial network architectures vary across cell types [28, 29, 30, 31], and transition between fragmented and hyperfused states in response to disease [32, 33, 34, 35, 36, 37], metabolic conditions [38, 39, 40], calcium signalling and apoptosis [41, 42], and cell cycle progression [43, 44]. These structural changes are thought to support important functions such as genetic complementation [45, 46], dilution of harmful reactive oxygen species (ROS) [47], power cabling [48, 49], quality control [50, 51, 52, 53], and modulation of mitochondrial heterogeneity across the population [54, 55]. These cellular functions rely on the connectivity-dependent diffusive spread of proteins, lipids, and/or ions through the underlying mitochondrial network [56, 21].
To elucidate how mitochondrial structure and dynamics govern the mixing of effectively diffusive particles, we introduce a quantitative framework for transport on spatially-embedded dynamic networks. We show that network connectivity governs spreading behavior by modulating several competing timescales: the cluster filling rate, encounter rate, and fusion rate, as well as the decay rate of the spreading material. In particular, we demonstrate a transition from rapid but low-dimensional spread in well-connected networks to slower three-dimensional spreading through an interacting population. Due to the different dimensionalities involved, particles with distinct diffusivities and networks with distinct structures can exhibit qualitatively different scaling behavior for dispersion from a source. We use the model to predict spreading rates for mitochondrial networks in three different human cell types, demonstrating the structural diversity and different regimes of relevance for these networks.
@@
We seek to quantify the rate of material dispersion through dynamic spatial networks ranging from the social to the physical regime. These limiting architectures, as well as a continuum of intermediate network connectivities, are encompassed by a simulation framework [57], illustrated in Fig. 1, consisting of $N_{0}$ interacting spherocylindrical units of length $\ell_{0}$, which move diffusively through a bounded spherical domain of radius $R$ with unit diffusivity $D_{1}$. Upon close encounters, pairs of units may fuse into larger clusters with rate constants $k_{u1}$ (for tip-tip fusions) or $k_{u2}$ (for tip-side fusions). Fission breaks connections at the nodes, with fission rate proportional to the number of attached edges ($k_{f}$ at degree-2 nodes and $1.5k_{f}$ at degree-3 nodes), in keeping with prior models [13]. At steady state, the system exhibits a characteristic mean cluster size $\langle n\rangle$ set by the balance of fusion and fission, and an associated cluster diffusivity $D_{n}$. Details of this structural model and parameterization are described in Ref. [57]; here we focus on the spreading of material throughout the resulting dynamic networks.
@@
We select at random a single unit, which serves as a particle source with fixed constant concentration $h_{0}=1/\ell_{0}$. Such a fixed source could represent, for instance, an import site for nucleus-encoded proteins [58], a site of regulated translation at a mitochondrial RNA granule or nucleoid [59], or a location of calcium entry at a contact with the endoplasmic reticulum [60]. The concentration field is propagated between connected units on the network with effective diffusivity $D_{p}$, using a finite volume approach [61] discretized at the level of individual units. Concentrations decay with a constant rate $k_{d}$ (representing removal from the network). The decay rate serves to set a relevant timescale, with the resulting calculations closely related to the question of how much material spreads from the source over a certain time. We quantify dispersion by computing the total amount of material ($S$) in the network at steady-state, excluding the source unit. Example steady-state snapshots are shown in Fig. 1b and Supplemental Video 1.
@@
Given the complex internal structure of mitochondria and the variety of functionally relevant biomolecules within them, particle diffusivities can vary broadly. For example, proteins in the inner mitochondrial membrane can be embedded in the extensive folds of cristae [62]. The cristae also serve as diffusive barriers to hinder transport of solutes in the matrix [63] and intermembrane space [64]. Our simplified modeling approach coarse-grains these complications into a single effective diffusivity $D_{p}$ representing the motion of a particle along the mitochondrial tubule axis. Reported effective diffusivities vary over orders of magnitude, from $\sim 0.004\mu\text{m}^{2}/\text{s}$ for ATP synthase components  [65], to $\sim 20\mu\text{m}^{2}/\text{s}$ for mitochondrial matrix proteins [66], with diffusivity of ions presumed even higher [67]. We consider a range of diffusivities $D_{p}=0.4-40\mu\text{m}^{2}/\text{s}$ in our simulations. The diffusivity of individual mitochondria is estimated to be much slower, at $D_{1}\approx 0.25\mu\text{m}^{2}/\text{min}$ [28, 68].
@@
The results below are reported in dimensionless units, relative to the length scale $2\ell_{0}$ and the timescale $1/k_{f}$. A reasonable estimate sets the mitochondrial unit length to $\ell_{0}=0.5\mu\text{m}$ and the fission timescale to $1/k_{f}\approx 2\text{ min}$ [57], consistent with experimental measurements of overall fission rate [69, 28]. The corresponding dimensionless diffusivities are $D_{p}=48-4800$ for the material, and $D_{1}=0.5$ for individual mitochondria.
Figure 2: Material spreading on static networks of different dimension $d$. Two different sets of networks are considered: near-linear structures (blue) with $k_{u2}=0.01k_{u1}$ and highly-branched structures (green) with $k_{u2}=10k_{u1}$. The simulations are run to steady state, and the network structures are then frozen. (a) Calculation of the network dimension for a single instance of each network type (shown in inset). The number of nodes within a given graph distance is plotted against the graph distance on log-log axes, with the slope giving the dimension, $d$. Averaging over 21 snapshots from 3 independent simulations yields effective dimensionalities $d=1.2,d=1.9$, respectively. (b) Total material delivered from a source unit on the static networks is plotted as a function of the diffusive lengthscale, $\lambda_{p}=\sqrt{D_{p}/k_{d}}$. Solid curves show exact solution (details in SI Appendix), averaged over replicate snapshots. Dashed-circle curves show approximation with the linear motif (inset i), applicable at short $\lambda_{p}$. Dashed-square curves show continuum solution on a fractal domain (inset ii), matching well at long $\lambda_{p}$.
@@ Results
In the limit where $\tau_{d}$ is the shortest timescale, material spreads primarily within a single connected component. This regime is relevant for highly fused networks with large cluster sizes. Such networks tend to be relatively static in their topology, with any fissions that occur rapidly followed by re-fusion with the same neighboring unit [57]. Recent evidence of collocalizing mitochondrial fusion and fission machinery at ER-mitochondria contact sites also points towards increased likelihood of re-fusion following a fission event [70].
For static networks, the steady-state distribution of material can be found by solving the diffusion equation on each one-dimensional edge, while matching boundary conditions at each node (details in SI Appendix). This method provides an exact solution for each individual network structure. However, to gain insight on the relevant scaling regimes, we turn to a mean-field continuum approach.
@@
We first consider the limit where the diffusive lengthscale $\lambda_{p}=\sqrt{D_{p}/k_{d}}$ is smaller or comparable to the length of linear segments between junctions. In the case of a completely linear snake-like network, the steady-state material content $S_{1}$ is given by the solution of the one-dimensional (1D) diffusion equation, $S_{1}=\frac{n_{b}\lambda_{p}}{\ell_{0}}\tanh\left(\frac{\ell_{0}\left<n\right>/n_{b}}{\lambda_{p}}\right)$ with $n_{b}=2$. When there are additional branches adjacent to the source ($d>1$), more material is able to enter the network. We thus define $n_{b}$ as the average number of edges directly connected to the source unit, and approximate the entire network cluster as a set of $n_{b}$ linear spokes, each of length $\ell_{0}\left<n\right>/n_{b}$, connected to a central source (Fig. 2iviv). $S_{1}$ then gives an estimate of the total material spreading over a short diffusive lengthscale $\lambda_{p}$ (Fig. 2, dashed-circle curves).
For the regime where material spreads over a more extensive network structure ($\lambda_{p}$ much greater than segment length), we consider diffusion through a $d$-dimensional medium. Here we approximate the spreading as spherically symmetric, with a $d$-sphere of radius $a_{n}$ (enclosing the source edge) maintained at fixed concentration, and a reflecting $d$-sphere of radius $R_{n}$ representing the outer boundary of the cluster. We set $R_{n}$ such that the average distance between two points in the $d$-dimensional domain is equal to the average graph distance between network nodes (details in SI Appendix). Assuming that the $\left<n\right$> network units are uniformly distributed within the continuum sphere, the radius $a_{n}$ is set proportionately to allow for 1 source unit within the inner sphere: $1/a_{n}^{d}=(\langle n\rangle-1)/(R_{n}^{d}-a_{n}^{d})$.
@@
Fig. 2 shows the correspondence between diffusion through a fractal continuum (dashed-square curves) and the exact solution (solid curves) for networks with different connectivities. The linear motif solution $S_{1}$ is a good estimate for small diffusive lengths, while the continuum solution $S_{d}$ is a better approximation for large $\lambda_{p}$. When the diffusive lengthscale is large enough ($\lambda_{p}>R_{n}$), the entire connected component saturates. Overall, material delivery is boosted by increased network branching [21], which increases both the network dimension and the network density (lower $R_{n}$ for the same number of network units).
@@
We next consider the regime of highly fragmented ‘social’ networks, where clusters are sufficiently small (and particle diffusivity sufficiently fast) that concentrations are fully equilibrated within each cluster: $\tau_{c}\ll\left\{\tau_{\text{d}},\tau_{\text{int}}\right\}$. In this limit, we can approximate the network as a system of $N=N_{0}/\left<n\right$> identical spherical units with effective radius $a$, each representing a cluster of uniform concentration. These effective units are capable of undergoing transient fusions whenever they are within contact radius $b$, with local rate constant $k_{u}$. Each fusion equilibrates the particle concentration in the two units involved, and is instantaneously followed by fission so that no larger structures are formed.
@@
$\frac{dh_{c}(t)}{dt}$ $=k_{u}\left[h_{0}\ell_{0}-h_{c}(t)\right]-\frac{I}{\rho v_{c}}-k_{d}h_{c}(t),$ (2b)
where $k_{u}\left[h_{0}\ell_{0}-h_{c}(t)\right]$ represents the injection of new material into the system via fusion with the source, $I=-4\pi b^{2}\rho D\left.\frac{dh(r,t)}{dr}\right|_{b}$ is the current of material leaving the contact zone, $\rho=(N_{0}-\langle n\rangle)/V$ is the density of network units, $V=4/3\pi(R^{3}-a^{3})$ is the domain volume, and $v_{c}=4/3\pi(b^{3}-a^{3})$ is the contact zone volume. Interactions between non-source units do not alter the mean-field concentrations, and the overall material in the system can only increase upon encounters with the source.
Figure 3: Material delivered from a source unit in a system of fragmented clusters. Solid lines show mean-field solutions (see Eq. 3) with unit cluster size $\langle n\rangle=1$, for different fusion rates $k_{u}$. Colored dots show explicit simulation results for a simplified system of interacting spheres with uniform size (inset). Dashed black lines show the limiting behavior for fusion-limited and encounter-limited regimes.
We compute the steady-state solution of Eq. 2 and integrate $h(r)$ over space to find the total amount of material ($S$) in the system, excluding the source unit (details in SI Appendix):
@@
The total material in the system exhibits distinct scaling behaviors (Fig. 3, dashed) depending on the relative timescales for decay, encounter ($\tau_{\text{enc}}=(4\pi Db\rho)^{-1}$), and fusion ($\tau_{u}=(k_{u}\rho v_{c})^{-1}$). In keeping with standard results for diffusion to a partially reactive target [73], the overall interaction time is given by the sum of waiting times for the two-step process: $\tau_{\text{int}}=\tau_{\text{enc}}+\tau_{u}$. In the limit of fast decay ($\tau_{d}\ll\tau_{\text{int}}$), the steady-state material in a social network of unit-size clusters approaches $S\rightarrow\rho v_{c}h_{c}=\rho v_{c}k_{u}/k_{d}$. In this limit, the total amount delivered is set by the number of units in the contact zone ($\rho v_{c}$) and how often they fuse with the source during the decay time ($k_{u}/k_{d}$), with no spatial spread.
@@
In a third limit, fusion of clusters is fast and decay is slow relative to the timescale of diffusive encounter. ($\tau_{u}\ll\tau_{\text{enc}}\ll\tau_{d}$). For this regime, the total material content is given by $S\rightarrow\rho zh_{c}\rightarrow 4\pi Db\rho/k_{d}$. This is an encounter-limited regime, where clusters arriving at the contact region fuse with the source nearly instantaneously ($h_{c}\approx h_{0}\ell_{0}=1$), and spreading is determined by the balance of arrival and decay rates.
Fig. 3 shows the full solution of the mean-field model (Eq. 3) for spreading on a fragmented network, with variable fusion rate $k_{u}$ and a unit cluster size $\left<n\right>=1$. The analytic calculations accurately reproduce simulations with diffusive spheres that exchange material via transient fusion (dots in Fig. 3, details in SI Appendix).
Figure 4: Spreading on simulated networks exhibits a transition between the static network and social network regime. The total material is plotted as a function of the decay time for different fusion rate constants $k_{u1}$ (solid lines). The analytic approximation (Eq. 4, dashed curves) encompases both the fractal continuum limit (for large clusters or fast decay) and the social network limit (for small clusters and slow decay). Plateaus at intermediate decay times correspond to filling of an individual cluster before interactions can occur. Insets show example simulation snapshots for the parameters indicated by the gray arrows. Parameters $N_{0}=250,\ell_{0}=0.5,R=5,k_{u2}=\frac{1}{3}k_{u1},k_{f}=1$ are used in both (a) and (b), with the particle diffusivity reduced from $D_{p}=4800$ in (a) to $D_{p}=48$ in (b).
@@
The model for spreading through a social network of interacting fragments can be expanded to approximate a regime with larger clusters. We use simulations of dynamic networks with different structures [57] to extract the effective parameters (mean cluster size $\left<n\right$>, effective unit diffusivity $D_{n}$, steric radius $a_{n}$, contact volume $v_{c}$, and effective local fusion rate $k_{u}$) for the simplified social network model (details in SI Appendix).
When individual clusters are homogeneously filled upon each interaction ($\tau_{c}<\left\{\tau_{d},\tau_{\text{int}},\tau_{f}\right\}$), the solution in Eq. 3 can be used directly with the appropriate values of the parameters $a,v_{c},k_{u}$. However, for slowly diffusing particles or very large clusters ($\tau_{c}>\tau_{d}$), the source cluster may be only partly full of material. The average concentration in the source cluster, $h_{0}^{(d)}$, is then determined by the static network solution (Eq. 1) within the cluster: $\ell_{0}h_{0}^{(d)}=(S_{d}(\sqrt{D_{p}/k_{d}})+1)/\langle n\rangle$. When another cluster fuses with the source cluster, fission may terminate the encounter before filling is complete ($\tau_{f}<\tau_{c}$). The amount transferred should then be scaled by the fraction of the new cluster that is filled prior to fission: $f=(S_{d}(\sqrt{D_{p}/k_{f}})+1)/\langle n\rangle$. These two corrections modify the dynamic equation for the concentration in the contact zone (Eq. 2b) and its solution (Eq. 3a,b) as follows: